OpenStudy (anonymous):

Miguel wants to use coordinate geometry to prove that the opposite sides of a rectangle are congruent. He places parallelogram ABCD in the coordinate plane so that A is (0, 0), B is (a, 0), C is (a, b), and D is (0, b). What formula can he use to determine the distance from point A to point B?

OpenStudy (anonymous):

\[A: (a-0)2 + (b-b)2 =a2 B: \square \root (a-0)2 + (0-0)2 =a2 =a C: (a-0)2 + (0-0) 2 = a2 D: \square \root (a-0)2 + (b-b)2 = a2 +a \]